OpenStudy (anonymous):

Given a line segment XZ of length 4. A point Y divides XZ such that \[\frac{ XY }{ YZ }=\frac{ 4 }{ \sqrt{32} }\] Find the lengths of XY and YZ.

hero (hero):

First, reduce the ratio. Then add the segment ratios together. Remember, XY + YZ = XZ

hero (hero):

\[\frac{4}{\sqrt{32}} = \frac{4}{\sqrt{(16)(2)}} = \frac{4}{4\sqrt{2}}= \frac{1}{\sqrt{2}} = \frac{\sqrt{2}}{2}\]

hero (hero):

So \[XY + YZ = \sqrt{2}x + 2x = 4\]

OpenStudy (anonymous):

So \[x=-2(\sqrt{2}-2)\] Then just plug that back in to the equation and I have my lengths.

OpenStudy (anonymous):

\[XY=4\sqrt{2}-4\] \[YZ=-4(\sqrt{2}-2)\] Thanks so much!